OpenStudy (anonymous):

Divide these polynomials. (4x4 + 10x3 + 12x2 + 5x - 6)/ (2x + 3) A. 2x3 - 2x2 + 3x - 1 B. 2x3 + 2x2 - 3x + 2 C. 2x3 + 2x2 + 3x - 2 D. 2x2 + 2x - 2

OpenStudy (anonymous):

option c is correct...

OpenStudy (anonymous):

\[\LARGE {4x^4+10x^3+12x^2+5x-6\over 2x+3}=\] \[\LARGE {(2x+3)(2x^3+2x^2+3x-2)\over 2x+3}=\] \[\LARGE {\cancel{(2x+3)}(2x^3+2x^2+3x-2)\over \cancel{2x+3}}=\] =Option C ;)